# MUDDFormer: Breaking Residual Bottlenecks in Transformers via Multiway Dynamic Dense Connections

### v1

@@ 1 Introduction
Dense connections (Huang et al., 2017) was proposed as a promising solution to the above issues, by allowing subsequent layers to directly access outputs of all preceding layers (Figure 2 (a)). It has shown effectiveness for CNNs with DenseNet (Huang et al., 2017), and recently for Transformers with DenseFormer (Pagliardini et al., 2024). However, these dense connection approaches use static (either fixed (Huang et al., 2017) or learnable (Pagliardini et al., 2024)) dense connection weights that are shared across sequence positions and different input streams of a Transformer block. As will be shown, this rigidity severely limits their expressive capacity in Transformers.
@@ 2.2 Dynamic Dense Connections
Overall, dynamic dense connections can be viewed as a form of depth-wise single-headed self-attention (Vaswani et al., 2017) with query and keys , and the attention weights are computed from the query side.
@@ 2.3 Multiway Dynamic Dense Connections
In a Transformer block, a single input is reused simultaneously as the query, key, value and residual of the MHA module (Figure 2 (e) top). These input streams play divergent roles and we hypothesize that they benefit from differentiated dense connectivity. To enable this, we first turn a normal Transformer block into a multi-input one by decoupling its input into four streams for query, key, value and residual, respectively (Eq. (7), Figure 2 (e) bottom), and then instantiate four DA modules, each specializing in one stream’s dense connectivity (Eq. (8), Figure 2 (d))222This is a logical view for clarity. In practice, these DSs can be combined for efficiency. See pseudocode in Appendix A.:
By making the dense connections multiway, the cross-layer communication bandwidth is further increased significantly. Multiway dynamic dense connections can be seen as depth-wise multi(4)-head attention. This vertical cross-layer attention can be composed with the horizontal cross-token attention in Transformer to form pathways adaptively, enhancing information flow across the whole model when performing in-context learning tasks.333For an illustrative example of this, see Figure 6. At this point, we finally obtain MUDDFormer by integrating static, dynamic and multiway dense connections. Complete pseudo-code for MUDDFormer is given in Appendix A.
@@ 2.5 Optional Normalization
It is similar to the hybrid-norm strategy used by recent models such as Gemma 2 (Team et al., 2024) and Grok-1 (xai org, 2024), though we apply it to DA modules instead of MHA/MLP modules. We use this PrePostDANorm variant to train the DeepNarrow models in scaling law experiments in Section 3.1 and the MUDDViT model in Appendix D.
@@ 2.6 Complexity Analysis
Table 1 shows the ratios of extra parameters and computation introduced by MUDD connections with both analytical results and typical concrete values. The derivations are in Appendix B. The ratio of extra parameters, i.e. parameter count of and of DA modules divided by that of the whole model, is proportional to the rectified depth/width ratio . The ratio of extra computation, i.e. FLOPs of generating MUDD connection weights and cross-layer aggregation divided by FLOPs of the whole forward pass, besides proportional to , decreases with . Both ratios are negligible for commonly used settings.
@@ 3 Experiments
Implementation Details We implement MUDDFormer model and training in JAX. We initialize the MUDD connection weight generating parameters and with and 0 respectively, and initialize the static weight vector with 1 at and 0 elsewhere. This reduces MUDDFormer to Transformer at the beginning of training, which is found to be critical for good performance. If PrePostDANorm is used, we initialize the scale parameters of Pre-DA and Post-DA RMSNorms with 1 and 1e-3, respectively, and initialize to 0 because is added as the residual after DA modules (Equation 10). For the other parameters outside DA modules, we use Xavier normal initializer.
Organization Our evaluation focuses on language modeling with decoder-only Transformer architecture, analyzing both pretraining scaling laws (Section 3.1) and downstream task performance (Section 3.2) with large scale training on the Pile dataset (Gao et al., 2020). Section 3.3 elucidates why MUDDFormer works through analyzing trained models, followed by efficiency analysis (Section 3.4) and ablations (Section 3.5). Extended vision experiments are provided in Appendix D.
@@ 3.1 Scaling Laws
Settings Table 2 (top half) specifies the model sizes and hyperparameters for scaling experiments, which are mostly taken from GPT-3 specifications (Brown et al., 2020). We untie input and output embedding matrices. We train with context length 2048 and set the number of training tokens to roughly match Chinchilla scaling laws (Hoffmann et al., 2022). The other hyperparameters are in Section C.1. In another set of scaling experiments we trade off some width for depth to train DeepNarrow models (Table 2 (bottom half)) to see if MUDD connections offer any benefits for this style of scaling.
@@
Baselines We compare MUDD with two recently proposed approaches to enhancing residual connections in Transformers: DenseFormer (Pagliardini et al., 2024) (same as the static dense connections described in Section 2.2) and Hyper-Connections (Zhu et al., 2025). We also compare to Transformer with dynamic dense connections (DDFormer) as described in Section 2.2. All these approaches are applied to an improved and now widely adopted Transformer architecture (Touvron et al., 2023) with rotary positional encoding (RoPE) (Su et al., 2024), SwiGLU MLP (Shazeer, 2020), etc. (often called Transformer++). We also include the plot for the original Transformer architecture used by GPT-3 as a comparison. The details for these baselines are in Section C.2.
Results Figure 3 plots Pile validation loss scaling curves of the models. While DenseFormer and Hyper-Connections show clear advantage over Transformer++, DDFormer outperforms these two baselines by adding dynamicity to dense connections. MUDDFormer further improves upon DDFormer by making the dense connections multiway, significantly outperforming all baselines on models ranging from 405M to 1.4B. It can be estimated that MUDDFormer-834M matches the loss of Transformer++ trained with 1.89× compute.
@@
3.2 Large Scaling Training and Downstream Evaluations
@@ 3.3 Analyzing and Understanding
Representation Collapse Figure 5 quantifies representation collapse through cosine similarity between the inputs of adjacent layers. While Pythia exhibits progressive collapse with 0.97 similarity in later layers, MDDDPythia maintains more distinct input representations, particularly in the value stream. Thanks to input stream decoupling and stream-specific aggregation, DA modules can freely aggregate distinct value input streams for MHA modules of each layer at each sequence position. These values will then be moved to other positions by MHA at this layer without polluting the residual stream of the current position. The relative importance of dense connections for the value stream is also evidenced by ablation studies (Section 3.5). We compare illustrative V-composition circuits444common and important in many tasks, e.g. Wang et al. (2023); Ni et al. (2025). For introduction, refer to Elhage et al. (2020). in Transformer and MUDDFormer in Figure 6 to highlight the benefit of MUDD connections and input stream decoupling, which results in more direct and cleaner information pathways. MUDD has similar effect on Q/K-composition circuits.
@@
Transformer models often exhibit null attention (Vig & Belinkov, 2019) where attention heads focus on the initial tokens or some special tokens as default “attention sinks” (Xiao et al., 2024) when no relevant tokens are found.
@@
MUDDPythia demonstrates ~2.4× higher activation ratio across layers, particularly in deeper layers.555Visualization of sampled attention patterns for heads in Pythia and MUDDPythia are shown in Appendix E.
@@ 3.4 Training and Inference Efficiency
Besides theoretical complexity analysis in Section 2.6, we assess training and inference efficiency of MUDDFormer compared with Transformer in real-world settings.
@@ 4 Related Work
Enhancing Residual Connections Despite the pervasive use of residual connections (He et al., 2016) in modern deep architectures, various approaches have been proposed to address their issues such as representational collapse and diminishing return for deeper models by strengthening cross-layer communication. Huang et al. (2017) introduced DenseNet for CNNs. Inspired by it, Pagliardini et al. (2024) proposed DenseFormer for Transformers, which uses Depth Weighted Averaging modules to aggregate outputs from all preceding layers with static, learnable weights. Most recently, Zhu et al. (2025) proposed Hyper-Connections (HC), an alternative to residual connections that uses both static and dynamic weights to adjust inter-layer dependencies. Other research has explored different forms of cross-layer attention (ElNokrashy et al., 2022; Fang et al., 2023; Wang et al., 2024) which retrieve or update representations across different layers in a more flexible manner. MUDDFormer is closely related to DenseFormer and HC but differs in critical ways. First, unlike DenseFormer, our MUDD connections dynamically compute per-position weights conditioned on the hidden states. Second, although HC uses a combination of static and dynamic weights to expand the hidden states, it does not employ explicit all-to-all cross-layer dense connectivity. Moreover, none of existing approaches consider decoupling the four input streams of a Transformer block by a multiway design, which is shown to bring significant performance gain in MUDDFormer.
@@
Cross-Layer KV Cache Optimization Brandon et al. (2024) proposes Cross-Layer Attention (CLA) to reduce Transformer KV cache size by sharing keys and values between adjacent layers, trading expressiveness for efficiency. Our MUDD connections enable cross-layer information flow between KV caches via dense connections on key and value streams. This enhances KV cache expressiveness and utility, improving in-context learning as evidenced by experiments. OmniNet (Tay et al., 2021a) achieves a fully global KV Cache receptive field by allowing each token to attend to all tokens in all layers. MUDDFormer attains a similar effect in a much more efficient way via composition of cross-layer dense connections and within-layer attention.
Intra-Layer Architectural Innovations Many other studies attempt to enhance the performance or efficiency of foundational sequence models with individual layers, including attention mechanisms (Ye et al., 2024; Leviathan et al., 2024; Liu et al., 2024), sub-quadratic linear attention or RNN/SSM architectures (Gu & Dao, 2024; Dao & Gu, 2024; Peng et al., 2024; Yang et al., 2024) and sparse Mixture-of-Experts (MoE) (Fedus et al., 2022; Dai et al., 2024). By contrast, MUDD connections focus on cross-layer communication, making it orthogonal and complementary to these approaches. We leave the exploration of combining MUDD connections with these within-layer optimizations for future work.
@@ Acknowledgements
Appendix A PyTorch Style Pseudo-code for MUDDFormer
@@
Appendix B Details of Complexity Analysis
Compared to Transformer++, extra compute and parameters in MUDDFormer are introduced by DA modules, increasing from bottom layer to top layer, due to varied hidden dim of DA at layer i. We approximately calculates and , the ratios of extra parameters and computation by considering an average layer in the middle of MUDDFormer. In this layer, . We assume in a typical Transformer architecture. We denote and . To simplify, we ignore RMSNorm because it is negible in terms of parameters and computation.
@@ Appendix B Details of Complexity Analysis
Appendix C Hyperparameters and Baselines for Scaling Law Experiments
C.1 Hyperparameters

### v2

@@ 1 Introduction
Dense connections (Huang et al., 2017) was proposed as a promising solution to the above issues, by allowing subsequent layers to directly access outputs of all preceding layers (Figure 2 (a)). It has shown effectiveness for CNNs with DenseNet (Huang et al., 2017), for encoder-decoder Transformers with Deep Transformers (Wang et al., 2019) and for decoder-only Transformers with DenseFormer (Pagliardini et al., 2024). However, these dense connection approaches use static (either fixed (Huang et al., 2017) or learnable (Wang et al., 2019; Pagliardini et al., 2024)) dense connection weights that are shared across sequence positions and different input streams of a Transformer block. As will be shown, this rigidity severely limits their expressive capacity in Transformers.
@@ 2.3 Multiway Dynamic Dense Connections
In a Transformer block, a single input is reused simultaneously as the query, key, value and residual of the MHA module (Figure 2 (e) top). These input streams play divergent roles and we hypothesize that they benefit from differentiated dense connectivity. To enable this, we first turn a normal Transformer block into a multi-input one by decoupling its input into four streams for query, key, value and residual, respectively (Eq. (7), Figure 2 (e) bottom), and then instantiate four DA modules, each specializing in one stream’s dense connectivity (Eq. (8), Figure 2 (d))333This is a logical view for clarity. In practice, these DSs can be combined for efficiency. See pseudocode in Appendix B.:
By making the dense connections multiway, the cross-layer communication bandwidth is further increased significantly. Multiway dynamic dense connections can be seen as depth-wise multi(4)-head attention. This vertical cross-layer attention can be composed with the horizontal cross-token attention in Transformer to form pathways adaptively, enhancing information flow across the whole model when performing in-context learning tasks.444For an illustrative example of this, see Figure 7. At this point, we finally obtain MUDDFormer by integrating static, dynamic and multiway dense connections. Complete pseudo-code for MUDDFormer is given in Appendix B.
@@ 2.5 Optional Normalization
It is similar to the hybrid-norm strategy used by recent models such as Gemma 2 (Team et al., 2024b) and Grok-1 (xai org, 2024), though we apply it to DA modules instead of MHA/MLP modules. We use this PrePostDANorm variant to train the DeepNarrow models in scaling law experiments in Section 3.1 and the MUDDViT model in Appendix F.
@@ 2.6 Complexity Analysis
Table 1 shows the ratios of extra parameters and computation introduced by MUDD connections with both analytical results and typical concrete values. The derivations are in Appendix C. The ratio of extra parameters, i.e. parameter count of and of DA modules divided by that of the whole model, is proportional to the rectified depth/width ratio . The ratio of extra computation, i.e. FLOPs of generating MUDD connection weights and cross-layer aggregation divided by FLOPs of the whole forward pass, besides proportional to , decreases with . Both ratios are negligible for commonly used settings.
@@ 3 Experiments
Implementation Details We implement MUDDFormer model and training in JAX. We initialize the MUDD connection weight generating parameters and with and 0 respectively, and initialize the static weight vector with 1 at and 0 elsewhere. This reduces MUDDFormer to Transformer at the beginning of training, which is found to be critical for good performance. If PrePostDANorm is used, we initialize the scale parameters of Pre-DA and Post-DA RMSNorms with 1 and 1e-3, respectively, and initialize to 0 because is added as the residual after DA modules (Eq. 10). For the other parameters outside DA modules, we use Xavier normal initializer.
Organization Our evaluation focuses on language modeling with decoder-only Transformer architecture, analyzing both pretraining loss scaling laws (Section 3.1 and Section 3.2) and downstream task performance (Section 3.3) with large scale training on the Pile dataset (Gao et al., 2020). Section 3.4 elucidates why MUDDFormer works through analyzing trained models, followed by efficiency analysis (Section 3.5) and ablations (Section 3.6). Extended vision experiments are provided in Appendix F.
@@ 3.1 Scaling Laws
Settings Table 2 (top half) specifies the model sizes and hyperparameters for scaling experiments, which are mostly taken from GPT-3 specifications (Brown et al., 2020). We untie input and output embedding matrices. We train with context length 2048 and set the number of training tokens to roughly match Chinchilla scaling laws (Hoffmann et al., 2022). The other hyperparameters are in Section D.1. In another set of scaling experiments we trade off some width for depth to train DeepNarrow models (Table 2 (bottom half)) to see if MUDD connections offer any benefits for this style of scaling.
@@
Baselines We compare MUDD with two recently proposed approaches to enhancing residual connections in Transformers: DenseFormer (Pagliardini et al., 2024) (same as the static dense connections described in Section 2.2) and Hyper-Connections (Zhu et al., 2025). We also compare to Transformer with dynamic dense connections (DDFormer) as described in Section 2.2. All these approaches are applied to an improved and now widely adopted Transformer architecture (Touvron et al., 2023) with rotary positional encoding (RoPE) (Su et al., 2024), SwiGLU MLP (Shazeer, 2020), etc. (often called Transformer++). We also include the plot for the original Transformer architecture used by GPT-3 as a comparison. The details for these baselines are in Section D.2.
Results Figure 3 plots Pile validation loss scaling curves of the models. While DenseFormer and Hyper-Connections show clear advantage over Transformer++, DDFormer outperforms them by adding dynamicity to dense connections. MUDDFormer further improves upon DDFormer by making the dense connections multiway, significantly outperforming all baselines on models ranging from 405M to 1.4B. It can be estimated that MUDDFormer-834M matches the loss of Transformer++ trained with 1.89× compute.
@@
3.2 MoE Models
Settings Transformer with Sparse Mixture-of-Experts (MoE) and MUDDFormer are both architectures with dynamic weights. MoE uses these weights to select experts within a layer while MUDDFormer uses the weights to aggregate outputs across layers. They are complementary approaches and can be combined. To empirically compare MoE and MUDDFormer, we train a Transforemr++ MoE model 1.8B-A405M with the same activated parameters as the 405M dense model, using the same training settings in the scaling law experiments. For MoE specific settings, we largely follow OLMoE (Muennighoff et al., 2025), using dropless token-choice routing to choose 2 experts out of 16 for each token with expert hidden dim 1408. The FLOPs of the MoE model is nearly identical to that of the dense model. We apply MUDD connections to both the dense and MoE models. All the models are trained on 26B tokens.
Results As shown in Figure 5, compared to the same Transformer++-405M, MUDDFormer-405M achieves ~50% loss reduction of Transformer++-1.8B-A405M MoE, which is 4.4× larger. MoE works by decoupling parameters and computation and relies on significantly expanding parameters for good performance, in contrast to MUDDFormer’s efficient utilization of both parameters and computation. More notably, MUDD demonstrates more benefits for the MoE model (loss 0.0641) than for the dense model (loss 0.0596), a result of particular practical significance given the increasing adoption of MoE architectures in frontier LLMs (Team et al., 2024a; Liu et al., 2024b; Team, 2025; Yang et al., 2025).
@@ 3.4 Analyzing and Understanding
Representation Collapse Figure 6 quantifies representation collapse through cosine similarity between the inputs of adjacent layers. While Pythia exhibits progressive collapse with 0.97 similarity in later layers, MDDDPythia maintains more distinct input representations, particularly in the value stream. Thanks to input stream decoupling and stream-specific aggregation, DA modules can freely aggregate distinct value input streams for MHA modules of each layer at each sequence position. These values will then be moved to other positions by MHA at this layer without polluting the residual stream of the current position. The relative importance of dense connections for the value stream is also evidenced by ablation studies (Section 3.6). We compare illustrative V-composition circuits555common and important in many tasks, e.g. Wang et al. (2023); Ni et al. (2025). For introduction, refer to Elhage et al. (2020). in Transformer and MUDDFormer in Figure 7 to highlight the benefit of MUDD connections and input stream decoupling, which results in more direct and cleaner information pathways. MUDD has similar effect on Q/K-composition circuits.
@@
Transformer models often exhibit null attention (Vig & Belinkov, 2019) where attention heads focus on the initial tokens or some special tokens as default “attention sinks” (Xiao et al., 2024b) when no relevant tokens are found.
@@
MUDDPythia demonstrates ~2.4× higher activation ratio across layers, particularly in deeper layers.666Visualization of sampled attention patterns for heads in Pythia and MUDDPythia are shown in Appendix G.
@@ 3.5 Training and Inference Efficiency
Besides theoretical complexity analysis in Section 2.6, we assess training and inference efficiency of MUDDFormer compared with Transformer in real-world settings.777For additional results of analytical and measured memory usage and training wall-clock time of MUDDFormer vs Transformer, see Appendix E.
@@ 4 Related Work
Enhancing Residual Connections Despite the pervasive use of residual connections (He et al., 2016) in modern deep architectures, various approaches have been proposed to address their issues such as representational collapse and diminishing return for deeper models by strengthening cross-layer communication. Huang et al. (2017) introduced densely connected convolutional networks (DenseNet) for image classification. Inspired by DenseNet, Pagliardini et al. (2024) proposed DenseFormer for Decoder-only Transformers, which uses Depth Weighted Averaging modules to aggregate outputs from all preceding layers with static, learnable weights. Similarly, Wang et al. (2019) proposed Dynamic Linear Combination of Layers (DLCL) in deep encoder-decoder Transformers for neural machine translation, which also used static and learnable dense connection weights. MUDD connections also have some linkage with HighwayNetworks (HN) (Srivastava et al., 2015). Both take inspiration from sequence model architectures and apply depthwise (HN from LSTM, MUDD from attention). HN is also the first to propose the critical concept of input dependent gating when mixing outputs between layers. Most recently, Zhu et al. (2025) proposed Hyper-Connections (HC), an alternative to residual connections that uses both static and dynamic weights to adjust inter-layer dependencies. Other research has explored different forms of cross-layer attention (ElNokrashy et al., 2022; Fang et al., 2023; Wang et al., 2024) which retrieve or update representations across different layers in a more flexible manner.
MUDDFormer is closely related to DenseFormer and HC but differs in critical ways. First, unlike DenseFormer, our MUDD connections dynamically compute per-position weights conditioned on the hidden states. Second, although HC uses a combination of static and dynamic weights to expand the hidden states, it does not employ explicit all-to-all cross-layer dense connectivity. Moreover, none of existing approaches consider decoupling the four input streams of a Transformer block by a multiway design, which is shown to bring significant performance gain in MUDDFormer.
@@
Cross-Layer KV Cache Optimization Brandon et al. (2024) proposes Cross-Layer Attention (CLA) to reduce Transformer KV cache size by sharing keys and values between adjacent layers, trading expressiveness for efficiency. Our MUDD connections enable cross-layer information flow between KV caches via dense connections on key and value streams. This enhances KV cache expressiveness and utility, improving in-context learning as evidenced by experiments. OmniNet (Tay et al., 2021a) achieves a fully global KV Cache receptive field by allowing each token to attend to all tokens in all layers. The authors reported that the additional omni (i.e. all-to-all) attention in OmniNet is computationally expensive and proposed to mitigate it by efficient attention variants. In contrast, MUDD connections is much more efficient because: 1) Only depth-wise aggregation (DA) is introduced, since the other omni attention paths can be formed by composition of DA and within-layer MHA; 2) DA is implemented as lightweight query-wise attention. The computation overhead, both theoretical and practical, is much lower.
Intra-Layer Architectural Innovations Many other studies attempt to enhance the performance or efficiency of foundational sequence models with individual layers, including attention mechanisms (Liu et al., 2024a; Xiao et al., 2024a; Ye et al., 2024; Leviathan et al., 2024), sub-quadratic linear attention or RNN/SSM architectures (Gu & Dao, 2024; Dao & Gu, 2024; Peng et al., 2024; Yang et al., 2024) and sparse Mixture-of-Experts (MoE) (Fedus et al., 2022; Dai et al., 2024). By contrast, MUDD connections focus on cross-layer communication, making it orthogonal and complementary to these approaches. We leave the exploration of combining MUDD connections with these within-layer optimizations for future work.
@@ Acknowledgements
Impact Statement
Our work introduces Multiway Dynamic Dense (MUDD) connections, a lightweight architectural innovation that fundamentally enhances Transformer-based foundation models. MUDD connections achieve up to 2.4× training efficiency gains while improving in-context learning capabilities, a critical requirement for real-world LLM applications. This breakthrough directly addresses the escalating computational and environmental costs of training ever-larger models, offering a sustainable pathway to high-performance AI without exponential parameter growth.
The open-source release of MUDDFormer architectures and pre-trained models will democratize access to state-of-the-art model efficiency techniques, particularly benefiting resource-constrained researchers and organizations. Enhanced attention mechanisms from MUDD’s multiway design could advance mechanistic interpretability research by producing more structured attention patterns and circuits.
While our method itself is architecture-focused, we acknowledge that more capable language models could potentially be misused for harmful content generation. However, MUDDFormer’s efficiency gains may paradoxically mitigate this risk by reducing the energy barrier to developing safer, smaller models that match larger counterparts’ performance. We commit to implementing rigorous model release protocols aligned with responsible AI practices.
Appendix A Dynamic Dense Connections as Depth-wise Self-Attention
Given a sequence , the output of dot-product self-attention at position is:
where ( is head dim) are input-dependent keys. Combining Eq. (5) and Eq. (6) in Section 2, the output of dynamic DA at layer for position is (operating depthwise across layers):
Comparing Eq. (11) and Eq. (12), plays the role of query projection () and ( is the inner dim of MLP, also the head dim) are keys as parameters independent of input. So dynamic DA can be seen as lightweight self-attention except:
Keys are independent of input;
A learnable positional bias is used;
Softmax is removed. Instead, GELU activation is applied to query (more like linear attention);
transformation is not used.
While theoretically these simplifications may impact the representation capacity, we empirically found that adding more sophisticated ingredients in DA (e.g. input dependent keys, softmax) does not bring improvement and slow down training.
Appendix B PyTorch Style Pseudo-code for MUDDFormer
@@
Appendix C Details of Complexity Analysis
Compared to Transformer++, extra compute and parameters in MUDDFormer are introduced by DA modules, increasing from bottom layer to top layer, due to varied hidden dim of DA at layer i. To estimate the overhead, we calculate and , the ratios of extra parameters and computation by analyzing an average middle layer in MUDDFormer. For this layer, the average hidden dimension of DA is . In a typical Transformer architecture we assume and define , . We omit RMSNorm because it is negligible in terms of parameters and computation.
@@ Appendix C Details of Complexity Analysis
In a Transformer++ of layers and model dimension, the number of parameters is ( for Attention and for FFN). In MUDDFormer, the layer adds parameters in and in of DA, so the ratio of extra parameters is as follows.
Approximating with and assuming :
Thus, the ratio of extra parameters scales linearly with the rectified depth/width ratio .
@@
In a Transformer++ model, the pretraining FLOPs is for a sequence of length . At layer of MUDDFormer, the extra FLOPs includes for generating dynamic dense weights (Eq. (6)) and for depthwise aggregate (Eq. (5)).
Similarly, we consider an average layer of MUDDFormer for simplification, then the extra FLOPS becomes .
Therefore, the ratio is approximately , decreasing with the rectified depth/width ratio.
Appendix D Hyperparameters and Baselines for Scaling Law Experiments
D.1 Hyperparameters